Compute the total quantity of heat (both in Kcal a Please show all work
Answer #1

Let the mass of the ice be m = 1.00 kg

We have to solve the problem stepwise

The data required for it is as follows

Specific heat of ice is Sice = 2.1 * 103 J/kg K

Latent heat of fusion of ice is Lice = 3.35 * 105 J/kg

Specific heat of water is Sw = 4.2 * 103 J/kg K

Latent heat of vaporization Lv = 2.26 * 106 J/kg

Specific heat of steam Ssteam= 2.08 * 103 J/kg K

1)Heat required to raise the temperature of from -100 C to 00 C

             Q1 = m * Sice * ∆T = 1 * 2.1 * 103 * ( 0 – ( -10 ) )

                   = 21000 J

2)Heat required to convert ice at 00 C to water at 00 C

                Q2 = m * Lice = 1 * 3.35 * 105 = 335000 J

3) Heat required to raise the temperature of water to 1000 C

                  Q3 = m * Sw * ∆T = 1 * 4.2 * 103 * ( 100 – 0 )

                        = 420000 J

4) Heat required to convert water at 1000 C to steam at 1000 C

                  Q4 = m * Lv = 1 * 2.26 * 106 = 2260000 J

5) Heat required to raise the temperature of stem from 1000 C to 1100 C

                 Q5 = m * Ssteam * ∆T = 1 * 2.08 * 103 * ( 110 – 100 )

                       = 20800 J

Now total heat supplied is Q = Q1 + Q2 + Q3 + Q4 + Q5

                                                      = 21000 + 335000 + 420000 + 2260000 + 20800

                                                     = 3056800 J

b) 1 Joule = 0.23 cal

so 3056800 J = 3056800 * 0.23 cal = 703064 cal = 703.064 kcal

So 703.064 kcal is supplied to convert ice to steam